OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3

OpenStudy (anonymous):

You have been given solutions to similar problems. Can you attempt the solution and let me know where you get stuck so I can help further?

OpenStudy (wasiqss):

make x the subject in 2nd eq, like x=3+y then subsitute this value of x in equation no 1

hero (hero):

y = x-3 y = (6-4x)/2 y = y 2(x-3) = 6-4x 2x - 6 = 6 - 4x 6x = 12 x = 12/6 x = 2

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

sorry sorry

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

hero's method is rite